# Fixed-domain curve counts for blow-ups of projective space

### v3

@@ Abstract.
We study the problem of counting pointed curves of fixed complex structure in blow-ups of projective space at general points. The geometric and virtual (Gromov-Witten) counts are found to agree asymptotically in the Fano (and some $(-K)$-nef) examples, but not in general. For toric blow-ups, geometric counts are expressed in terms of integrals on products of Jacobians and symmetric products of the domain curves, and evaluated explicitly in genus 0 and in the case of ${\mathsf{Bl}}_{q}(\mathbb{P}^{r})$. Virtual counts for ${\mathsf{Bl}}_{q}(\mathbb{P}^{r})$ are also computed via the quantum cohomology ring.
@@ 1.1. Curve-counting with fixed domain
Let $X$ be a non-singular, irreducible, projective variety over $\mathbb{C}$ of dimension $r$. Curve-counting problems on $X$ are traditionally formulated in Gromov-Witten theory as intersection numbers on the space of stable maps ¯$\overline{\mathcal{M}}_{g,n}(X,\beta)$ against the virtual fundamental class
[¯$[\overline{\mathcal{M}}_{g,n}(X,\beta)]^{\mathrm{vir}}\in A_{\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))}(\overline{\mathcal{M}}_{g,n}(X,\beta)).$
@@
We will be concerned with such problems where, in addition, the complex structure of the domain curve $(C,p_{1},\ldots,p_{n})$ is fixed (and general). Gromov-Witten counts are therefore obtained by additionally restricting to the pullback of a point under the forgetful map ¯$\overline{\mathcal{M}}_{g,n}(X,\beta)\to\overline{\mathcal{M}}_{g,n}$.
There are two salient features of the fixed-domain version of the problem: first, degenerating $(C,p_{1},\ldots,p_{n})$ to a union of rational curves, the answers may be expressed purely in terms of the small quantum cohomology ring of $X$ [5, Theorem 1.3], and therefore tend to be simpler than arbitrary Gromov-Witten invariants. Second, the restriction to a general curve often avoids the most pathological subvarieties of ¯$\overline{\mathcal{M}}_{g,n}(X,\beta)$, and therefore, the fixed-domain Gromov-Witten invariants are much more often enumerative [17, 3]. Furthermore, even when enumerativity fails, the corresponding geometric count, which is different, is usually well-defined and may still be accessible via more subtle methods.
@@
Fix integers $g,n\geq 0$ satisfying the stability condition $02g-2+n>$ 0 so that the moduli space ¯$\overline{\mathcal{M}}_{g,n}$ of $n$-pointed, genus $g$ stable curves is well-defined. Fix $\beta\in H_{2}(X,\mathbb{Z})$ an effective curve class satisfying the condition
@@
Let ¯$\overline{\mathcal{M}}_{g,n}(X,\beta)$ be the moduli space of genus $g$, $n$-pointed stable maps to $X$ in class $\beta$ and assume that the dimensional constraint
@@
The geometric Tevelev degree $\mathsf{Tev}^{X}_{g,n,\beta}\in\mathbb{Z}$ of $X$ is defined in [17, Definition 2] under the further assumption that the restriction of $\tau$ to maps with smooth domain
@@
has reduced and 00-dimensional general fiber, in which case its cardinality is by definition equal to $\mathsf{Tev}^{X}_{g,n,\beta}$. The integrality of $\mathsf{Tev}^{X}_{g,n,\beta}$ follows from the absence of automorphisms in a general fiber of $\tau$, see [17, Remark 4].
We record here the following crucial lemma, see [17, Proposition 14, proof of Proposition 22]. It does not require the assumption (1).
Lemma 1.
@@
More generally, let $Z\subset\overline{\mathcal{M}}_{g,n}(X,\beta)$ be an irreducible component whose general point $f:C\to X$ has the property that $C$ is a union of a smooth genus $g$ curve $C_{{\text{sp}}}$ containing $m$ of the marked points, and $n-m$ rational tails attached to $C_{{\text{sp}}}$, each containing a marked point. Suppose that $Z$ dominates ${\mathcal{M}}_{g,n}\times X^{n}$ and that $m\geq g+1$. Then, $Z$ is generically smooth of dimension equal to the virtual dimension.
In particular, $\mathsf{Tev}^{X}_{g,n,\beta}$ is well-defined under assumption (1) whenever $n\geq g+1$. This bound is not sharp: for example, when $X=\mathbb{P}^{r}$, the geometric Tevelev degrees are also well-defined whenever $n\geq r+1$ (in which case $f:C\to\mathbb{P}^{r}$ is necessarily non-degenerate) by Brill-Noether theory. It is an interesting question to determine for other $X$ whether the bound $n\geq g+1$ can be improved.
Three questions emerge:
@@ 1.2. Curve counts on $\mathbb{P}^{r}$
In fact, this formula had been obtained much earlier [4] before the modern theory of the Gromov-Witten virtual fundamental class was available. It is also a consequence of the Vafa-Intrilligator formula [20, 19].
The virtual and geometric counts agree when $d$ is large compared to $r$ and $g$:
Theorem 3.
@@
Obtaining geometric counts in low degree for $\mathbb{P}^{r}$ is much more difficult; the virtual and geometric counts no longer agree. For $\mathbb{P}^{1}$, we have:
Theorem 4.
@@
Binomial coefficients $\binom{g}{j}$ with $j<0$ are interpreted to vanish, recovering Theorem 3 when $d\geq r+1$. See [11, §2] for notation for the Schubert cycles appearing in the last formula.
@@ 1.3.1. Enumerativity
Fixed-domain virtual curve counts are much more likely to be enumerative than arbitrary Gromov-Witten invariants. For example, we have already seen that virtual and geometric counts for $\mathbb{P}^{r}$ agree in large degree, whereas higher-genus Gromov-Witten invariants of $\mathbb{P}^{r}$ typically fail to be enumerative. This phenomenon was first studied systematically in [17].
We assume throughout that the dimensional constraint (1) holds.
Definition 5.
@@
    For all $g\geq 0$, there exists a constant $C(X,g)$ (depending only on $X$ and $g$, but not on $\beta$) for which, if $n\geq C(X,g)$, then the general fiber of the map (2)
@@
If $\tau$ has general fiber contained in ${\mathcal{M}}_{g,n}(X,\beta)$ and furthermore $n\geq g+1$, then by Lemma 1, the general fiber of $\tau$ is reduced and 00-dimensional, so $\mathsf{Tev}^{X}_{g,n,\beta}=\mathsf{vTev}^{X}_{g,n,\beta}$, i.e. $\mathsf{vTev}^{X}_{g,n,\beta}$ is enumerative. Thus, we have:
Proposition 6.
@@
In [17], SAE is proven for homogeneous spaces for linear algebraic groups (see [17, Theorem 10]) and non-singular hypersurfaces of very low degree (see [17, Theorem 11, Corollary 34]), implying the conclusion of Proposition 6 in these examples. As it is difficult to imagine how one could obtain this conclusion for $X$ without SAE, the following question is natural.
Question 7.
@@
It was originally expected that all Fano $X$ satisfy SAE [17, Speculation 12 and §4]. However, several families of counterexamples to this speculation have been constructed in [3, §3]. The SAE property therefore seems somewhat more subtle.
Our first main set of results concerns SAE for $X$ given by the blow-up of $\mathbb{P}^{r}$ at distinct points $q_{1},\ldots,q_{\ell}$. We have the following negative result:
Theorem 8.
@@
Such $X$ are not Fano, and indeed, the existence of negative rational curves is used in an in essential way. In fact, we establish a more general criterion for failure of SAE which also applies to various special configurations of points $q_{j}$ (e.g., where some three of the $q_{j}$ are collinear), see Proposition 15.
@@
    (Theorem 24) $X$ is del Pezzo, i.e. $r=2$, and the $\ell\leq 8$ points satisfy the property that no three lie on a line, no six lie on a conic, and, if $\ell=8$, the points do not all lie on a cubic singular at one of the $q_{j}$.
@@
    (Theorem 32) $r=3,\ell\leq 4$, and the $q_{j}$ are (linearly) general.
@@
    (Theorem 23) $r$ is arbitrary and $\ell=1$.
In Theorem 9, examples 1 and 3 are Fano; in fact, these are all examples of Fano varieties obtained by blowing up $\mathbb{P}^{r}$ at general points. In example 2, $X$ is merely $(-K_{X})$-nef when $\ell\geq 2$; these are the first examples of non-Fano varieties proven to satisfy SAE.
@@ 1.3.2. Geometric calculations
Let $\mathsf{H}\in H^{2}(X)$ denote the pullback of the hyperplane class from $\mathbb{P}^{r}$, and let $\mathsf{E}_{1},\ldots,\mathsf{E}_{\ell}\in H^{2}(X)$ be the classes of the exceptional divisors. Let $\mathsf{H}^{\vee},\mathsf{E}^{\vee}_{1},\ldots,E^{\vee}_{\ell}\in H_{2}(X)$ be the corresponding dual basis, and let
@@
Setup Write $S:=\mathsf{Jac^{d}}(C)\times\mathsf{Sym}^{k_{1}}(C)\times\ldots\times\mathsf{Sym}^{k_{\ell}}(C)$. Over $S\times C$, we have the following:
@@ 2. $\bullet$
(3) $d-\sum_{i\in I}k_{i}>2g-1\text{ for all }I\subseteq\{1,\ldots,\ell\}\text{ such that }|I|\leq r,$
in order to ensure that $\mathcal{E}$ is a vector bundle. We will see later in §3.1 that we will also want this to remain true upon replacing $\mathcal{L}$ with $\mathcal{L}(-p_{i})$, hence the appearance of $12g-$ 1 rather than $22g-$ 2 on the right hand side.
@@
Then, we have:
Proposition 10 (Integral formula).
@@
The condition $n-d\geq g+1$ is needed to rule out extraneous contributions when the sections $f_{j}$ are identically zero (see Lemma 40 and Remark 41), and also ensure dimensional transversality later (Lemma 50). The factor $\widetilde{\mathsf{H}}^{r}+\sigma_{1}\widetilde{\mathsf{H}}^{r-1}+\cdots+\sigma_{r}$ is the class of a subscheme $V(x_{i})$ of $\mathbb{P}$ corresponding to the condition that $f(p_{i})=x_{i}$. However, care is needed to deal with the possibility that the sections $f_{j}$ vanish simultaneously at $p_{i}$; a precise definition of $V(x_{i})$ is given in §3.1. The computation of the class $[V(x_{i})]$ and the fact that the $V(x_{i})$ intersect transversely at points enumerated by $\mathsf{Tev}_{g,n,\beta}^{X}$ are both subtle.
Pushing forward to $S=\mathsf{Jac^{d}}(C)\times\mathsf{Sym}^{k_{1}}(C)\times\cdots\times\mathsf{Sym}^{k_{\ell}}(C)$ and using Grothendieck-Riemann-Roch, we obtain (see §3.2.1 for notation):
Theorem 11.
@@
We extract from Theorem 11 relatively simple formulas in two special cases.
Theorem 12 ($g=0$ specialization).
@@
It is often the case that the formula above gives zero. For example, take $\ell=1$ and $\beta=d\mathsf{H}^{\vee}+k\mathsf{E}^{\vee}$ (where we write simply $\mathsf{E}_{1}^{\vee}=\mathsf{E}^{\vee}$), where $(r-1)k\leq d$. Assuming (1), the inequalities (3), and (5) are satisfied. Then, Theorem 12 reads simply
@@
Specializing Theorem 11 to $\ell=1$ in arbitrary genus yields:
Theorem 13 ($\ell=1$ specialization).
@@
Note that all of our geometric results require the inequalities (3) and (5). If $g$ and $k_{j}$ are viewed as fixed, then both inequalities hold for sufficiently large $d$, so our results may be viewed as “large degree” counts in analogy with Theorem 3. The low degree counts seem, unsurprisingly, more difficult, and are left open.
@@ 1.3.3. Virtual calculations
Via a computation in the quantum cohomology ring $QH^{*}(X)$, we also obtain:
Theorem 14.
@@
Thus, the formula matches that of Theorem 13, but holds in a wider range.

### v4

@@ Abstract.
We study the problem of counting pointed curves of fixed complex structure in blow-ups of projective space at general points. The geometric and virtual (Gromov–Witten) counts are found to agree asymptotically in the Fano (and some (K)(-K)-nef) examples, but not in general. For toric blow-ups, geometric counts are expressed in terms of integrals on products of Jacobians and symmetric products of the domain curves, and evaluated explicitly in genus 0 and in the case of ${\mathsf{Bl}}_{q}(\mathbb{P}^{r})$. Virtual counts for ${\mathsf{Bl}}_{q}(\mathbb{P}^{r})$ are also computed via the quantum cohomology ring.
@@ 1.1. Curve-counting with fixed domain
Let $X$ be a non-singular, irreducible, projective variety over $\mathbb{C}$ of dimension $r$. Curve-counting problems on $X$ are traditionally formulated in Gromov–Witten theory as intersection numbers on the space of stable maps $\overline{\mathcal{M}}_{g,n}(X,\beta)$ against the virtual fundamental class
$\overline{\mathcal{M}}_{g,n}(X,\beta)]^{\mathrm{vir}}\in A_{\mathrm{vdim}(\overline{\mathcal{M}}_{g,n}(X,\beta))}(\overline{\mathcal{M}}_{g,n}(X,\beta)).$
@@
We will be concerned with such problems where, in addition, the complex structure of the domain curve $(C,p_{1},\ldots,p_{n})$ is fixed (and general). Gromov–Witten counts are therefore obtained by additionally restricting to the pullback of a point under the forgetful map $\overline{\mathcal{M}}_{g,n}(X,\beta)\to\overline{\mathcal{M}}_{g,n}$.
There are two salient features of the fixed-domain version of the problem: first, degenerating $(C,p_{1},\ldots,p_{n})$ to a union of rational curves, the answers may be expressed purely in terms of the small quantum cohomology ring of $X$ [5, Theorem 1.3], and therefore tend to be simpler than arbitrary Gromov–Witten invariants. Second, the restriction to a general curve often avoids the most pathological subvarieties of $\overline{\mathcal{M}}_{g,n}(X,\beta)$, and therefore, the fixed-domain Gromov–Witten invariants are much more often enumerative [16, 3]. Furthermore, even when enumerativity fails, the corresponding geometric count, which is different, is usually well-defined and may still be accessible via more subtle methods.
@@
Fix integers $g,n\geq 0$ satisfying the stability condition 2g2+n>02g-2+n>0 so that the moduli space $\overline{\mathcal{M}}_{g,n}$ of $n$-pointed, genus $g$ stable curves is well-defined. Fix $\beta\in H_{2}(X,\mathbb{Z})$ an effective curve class satisfying the condition
@@
Let $\overline{\mathcal{M}}_{g,n}(X,\beta)$ be the moduli space of genus $g$, $n$-pointed stable maps to $X$ in class $\beta$ and assume that the dimensional constraint
@@
The geometric Tevelev degree $\mathsf{Tev}^{X}_{g,n,\beta}\in\mathbb{Z}$ of $X$ is defined in [16, Definition 2] under the further assumption that the restriction of $\tau$ to maps with smooth domain
@@
has reduced and 0-dimensional general fiber, in which case its cardinality is by definition equal to $\mathsf{Tev}^{X}_{g,n,\beta}$. The integrality of $\mathsf{Tev}^{X}_{g,n,\beta}$ follows from the absence of automorphisms in a general fiber of $\tau$, see [16, Remark 4].
We record here the following crucial lemma, see [16, Proposition 14, proof of Proposition 22]. It does not require the assumption (1).
Lemma 1.1.
@@
More generally, let $Z\subset\overline{\mathcal{M}}_{g,n}(X,\beta)$ be an irreducible component whose general point $f:C\to X$ has the property that $C$ is a union of a smooth genus $g$ curve $C_{{\text{sp}}}$ containing $m$ of the marked points, and nmn-m rational tails attached to $C_{{\text{sp}}}$, each containing a marked point. Suppose that $Z$ dominates ${\mathcal{M}}_{g,n}\times X^{n}$ and that $m\geq g+1$. Then, $Z$ is generically smooth of dimension equal to the virtual dimension.
In particular, $\mathsf{Tev}^{X}_{g,n,\beta}$ is well-defined under assumption (1) whenever $n\geq g+1$. This bound is not sharp: for example, when $X=\mathbb{P}^{r}$, the geometric Tevelev degrees are also well-defined whenever $n\geq r+1$ (in which case $f:C\to\mathbb{P}^{r}$ is necessarily non-degenerate) by Brill–Noether theory. It is an interesting question to determine for other $X$ whether the bound $n\geq g+1$ can be improved.
Three questions emerge.
@@ 1.2. Curve counts on $\mathbb{P}^{r}$
In fact, this formula had been obtained much earlier [4] before the modern theory of the Gromov–Witten virtual fundamental class was available. It is also a consequence of the Vafa–Intriligator formula [21, 20].
The virtual and geometric counts agree when $d$ is large compared to $r$ and $g$.
Theorem 1.3.
@@
Obtaining geometric counts in low degree for $\mathbb{P}^{r}$ is much more difficult; the virtual and geometric counts no longer agree. For $\mathbb{P}^{1}$, we have the following.
Theorem 1.4.
@@
Binomial coefficients $\binom{g}{j}$ with $j<0$ are interpreted to vanish, recovering Theorem 1.3 when $d\geq r+1$. See [12, §2] for notation for the Schubert cycles appearing in the last formula.
@@ 1.3.1. Enumerativity
Fixed-domain virtual curve counts are much more likely to be enumerative than arbitrary Gromov–Witten invariants. For example, we have already seen that virtual and geometric counts for $\mathbb{P}^{r}$ agree in large degree, whereas higher-genus Gromov–Witten invariants of $\mathbb{P}^{r}$ typically fail to be enumerative. This phenomenon was first studied systematically in [16].
Definition 1.5.
@@
    For all $g\geq 0$, there exists a constant $C(X,g)$ (depending only on $X$ and $g$, but not on $\beta$) for which, if $n\geq C(X,g)$ and the dimensional constraint (1) is satisfied, then the general fiber of the map (2)
@@
Recall that we require 2g2+n>02g-2+n>0, so when $g=0$, it must be the case that $n\geq 3$. Thus, (2) above asserts that when $g=0$, the condition on the general fiber holds for all values of $n$ that we consider. If $\tau$ has general fiber contained in ${\mathcal{M}}_{g,n}(X,\beta)$ and furthermore $n\geq g+1$, then by Lemma 1.1, the general fiber of $\tau$ is reduced and 0-dimensional, so $\mathsf{Tev}^{X}_{g,n,\beta}=\mathsf{vTev}^{X}_{g,n,\beta}$, i.e. $\mathsf{vTev}^{X}_{g,n,\beta}$ is enumerative. Thus, we have the following.
Proposition 1.6.
@@
In [16], SAE is proven for homogeneous spaces for linear algebraic groups (see [16, Theorem 10]) and non-singular hypersurfaces of very low degree (see [16, Theorem 11, Corollary 34]), implying the conclusion of Proposition 1.6 in these examples. As it is difficult to imagine how one could obtain this conclusion for $X$ without SAE, the following question is natural.
Question 1.7.
@@
It was originally expected that all Fano $X$ satisfy SAE [16, Speculation 12 and §4]. However, several families of counterexamples to this speculation have been constructed in [3, §3]. A symplectic analog of [16, Speculation 12] has more recently been established in [6].
Our first main set of results concerns SAE for $X$ given by the blow-up of $\mathbb{P}^{r}$ at distinct points $q_{1},\ldots,q_{\ell}$. We have the following negative result.
Theorem 1.8.
@@
Such $X$ are not Fano, and indeed, the existence of negative rational curves is used in an in essential way. In fact, we establish a more general criterion for failure of SAE which also applies to various special configurations of points $q_{j}$ (e.g., where some three of the $q_{j}$ are collinear), see Proposition 2.1.
@@
    (Theorem 2.9) $X$ is del Pezzo, i.e. $r=2$, and the $\ell\leq 8$ points satisfy the property that no three lie on a line, no six lie on a conic, and, if $\ell=8$, the points do not all lie on a cubic singular at one of the $q_{j}$.
@@
    (Theorem 2.17) $r=3,\ell\leq 4$, and the $q_{j}$ are (linearly) general.
@@
    (Theorem 2.8) $r$ is arbitrary and $\ell=1$.
In Theorem 1.9, examples 1 and 3 are Fano; in fact, these are all examples of Fano varieties obtained by blowing up $\mathbb{P}^{r}$ at general points. In example 2, $X$ is merely $(-K_{X})$-nef when $\ell\geq 2$; these are the first examples of non-Fano varieties proven to satisfy SAE.
@@ 1.3.2. Geometric calculations
Let $\mathsf{H}\in H^{2}(X)$ denote the pullback of the hyperplane class from $\mathbb{P}^{r}$, and let $\mathsf{E}_{1},\ldots,\mathsf{E}_{\ell}\in H^{2}(X)$ be the classes of the exceptional divisors. Let $\mathsf{H}^{\vee},\mathsf{E}^{\vee}_{1},\ldots,\mathsf{E}^{\vee}_{\ell}\in H_{2}(X)$ be the corresponding dual basis, and let
@@
Setup Write $S:=\mathsf{Jac^{d}}(C)\times\mathsf{Sym}^{k_{1}}(C)\times\ldots\times\mathsf{Sym}^{k_{\ell}}(C)$. Over $S\times C$, we have the following.
@@ 2. $\bullet$
[…]
in order to ensure that $\mathcal{E}$ is a vector bundle. We will see later in §3.1 that we will also want this to remain true upon replacing $\mathcal{L}$ with $\mathcal{L}(-p_{i})$, hence the appearance of 2g12g-1 rather than 2g22g-2 on the right hand side.
@@
Then, we have the following.
Proposition 1.10 (Integral formula).
@@
The condition $n-d\geq g+1$ is needed to rule out extraneous contributions when the sections $f_{j}$ are identically zero (see Lemma 3.8 and Remark 3.9), and also ensure dimensional transversality later (Lemma 3.18). The factor $\widetilde{\mathsf{H}}^{r}+\sigma_{1}\widetilde{\mathsf{H}}^{r-1}+\cdots+\sigma_{r}$ is the class of a subscheme $V(x_{i})$ of $\mathbb{P}$ corresponding to the condition that $f(p_{i})=x_{i}$. However, care is needed to deal with the possibility that the sections $f_{j}$ vanish simultaneously at $p_{i}$; a precise definition of $V(x_{i})$ is given in §3.1. The computation of the class $V(x_{i})]$ and the fact that the $V(x_{i})$ intersect transversely at points enumerated by $\mathsf{Tev}_{g,n,\beta}^{X}$ are both subtle.
Pushing forward to $S=\mathsf{Jac^{d}}(C)\times\mathsf{Sym}^{k_{1}}(C)\times\cdots\times\mathsf{Sym}^{k_{\ell}}(C)$ and using Grothendieck-Riemann-Roch, we obtain (see §3.2.1 for notation) the following.
Theorem 1.11.
@@
We extract from Theorem 1.11 relatively simple formulas in two special cases.
Theorem 1.12 ($g=0$ specialization).
@@
It is often the case that the formula above gives zero. For example, take $\ell=1$ and $\beta=d\mathsf{H}^{\vee}+k\mathsf{E}^{\vee}$ (where we write simply $\mathsf{E}_{1}^{\vee}=\mathsf{E}^{\vee}$), where $(r-1)k\leq d$. Assuming (1), the inequalities (3), and (5) are satisfied. Then, Theorem 1.12 reads simply
@@
Specializing Theorem 1.11 to $\ell=1$ in arbitrary genus yields the following.
Theorem 1.13 ($\ell=1$ specialization).
@@
Note that all of our geometric results require the inequalities (3) and (5). If $g$ and $k_{j}$ are viewed as fixed, then both inequalities hold for sufficiently large $d$, so our results may be viewed as “large degree” counts in analogy with Theorem 1.3. The low degree counts seem, unsurprisingly, more difficult, and are left open.
@@ 1.3.3. Virtual calculations
Via a computation in the quantum cohomology ring $QH^{*}(X)$, we also obtain the following.
Theorem 1.14.
@@
Thus, the formula matches that of Theorem 1.13, but holds in a wider range.